# Retro-Expert: Collaborative Reasoning for Interpretable Retrosynthesis

### v1

@@ Introduction
In this objective function, $\pi_{\theta}$ represents the policy model being optimized, while $\pi_{ref}$ is a reference model used to regularize the policy update, with their divergence measured by $\mathbb{D}_{kl}$. The term $q$ denotes a query sampled from the dataset, comprising the candidate set $\mathcal{T}$ and any external domain knowledge $k$. The output $y$ generated by the policy model includes both the complete reasoning path $\mathcal{T}_{LLM}$ and the final predicted reactants $a$. The optimization aims to maximize the expected reward $r_{\phi}(q,y)$. To effectively optimize this objective, we adopt the Group Relative Policy Optimization (GRPO) algorithm (Shao et al. 2024). Furthermore, we design a novel rule-based, multi-stage reward mechanism tailored to enhance the LLM’s reasoning over the candidate outputs from specialized retrosynthesis models.
While previous methods focus their reward signals solely on the correctness of the final reactant, we emphasize that the logical soundness of the reasoning path is equally critical. In practice, a chemist’s trust is earned not by a single, opaque prediction, but by an interpretable reasoning process that aligns with established chemical principles. To this end, the KGPO framework utilizes a composite reward function that evaluates both the process and the outcome. This function is specifically designed to leverage domain knowledge from specialized models, guiding the LLM’s reasoning toward chemical validity while maintaining high accuracy in the final prediction. For a sampled decision path $y$, it is defined as:
where the term $r_{i}$ represents the binary stage reward for each of the $n$ sub-tasks, assigned a value of 1 if the intermediate prediction is correct for that stage and 0 otherwise. $r_{\text{reactant}}$ is the reward for the correctness of the final predicted reactants and $r_{\text{format}}$ aim to ensure the required output format. The $\alpha$ terms represent the weights of the corresponding reward.
@@
Experiments
Experimental Setup
Dataset.
@@
We conduct experiments using the benchmark dataset USPTO-50K (Schneider, Stiefl, and Landrum 2016) which contains 50,000 atom-mapped reaction records, and adopt the same training/validation/test splits in 8:1:1 as prior works (Dai et al. 2019). Based on previous methods (Somnath et al. 2021), we canonicalize the product SMILES and reassign the atom-mapping to the corresponding reactant SMILES based on the canonical ordering.
Baselines.
@@
We compare Retro-Expert against various strong baselines, categorized into two primary classes: (1) specialized models based approaches, encompassing template-based (LocalRetro (Chen and Jung 2021), GLN (Dai et al. 2019)), semi-template-based (GraphRetro (Somnath et al. 2021), RetroPrime (Wang et al. 2021), Graph2Edits (Zhong et al. 2023)), and template-free (Retroformer (Yao et al. 2023), UAlign (Zeng et al. 2024)) models; (2) LLM-based methods, including chemistry-specialized LLMs and general LLMs. To ensure a fair comparison, we follow each model’s default evaluation settings and provide necessary predictions from specialized models (e.g., reaction type classifiers) to support the retrosynthesis prediction.
Evaluation Metrics.
@@
Inspired by (Zhao et al. 2024), we adopt two sets of metrics to evaluate the performance between the predicted and golden standardized SMILES. The first set focuses on the textual similarity, using Top-1 accuracy, BLEU, and levenshtein distance. The second set evaluates the chemical similarity, encompassing the validity of the generated SMILES and fingerprint tanimoto similarity (i.e., MACCS, RDK, Morgan). Additionally, we compare the performance of Retro-Expert against retrosynthetic specialized models using Top-1 accuracy.
Implementation Details.
@@
We employ three specialized models for expert-level retrosynthesis analysis: T5Chem (Lu and Zhang 2022) for reaction type prediction, and GraphRetro (Somnath et al. 2021) for both reaction center localization and reactant prediction. During inference, ANY specialized models can be used to provide candidate results for sub-tasks. To balance accuracy and optimization efficiency, we use Top-3 candidate predictions from each model. We utilize Qwen2.5-7B-Instruct as the LLM and train it with reinforcement learning using 9k samples. During training, to prevent reward hacking, we distribute the correct predictions from the specialized model among the top three candidate positions in a 5:3:2 ratio. This strategy prevents the model from developing a positional bias (e.g., always selecting the first candidate) and forces it to learn the intrinsic quality of predictions.
Results & Discussion
Superiority over LLM-based Approaches. As detailed in Table 1, Retro-Expert significantly outperforms both its underlying base model (Qwen2.5-7B-Instruct) and other chemistry-focused foundation models, achieving a Top-1 Accuracy improvement of over 22.59%. This demonstrates that our knowledge-guided RL reasoning is substantially more effective than the standard SFT paradigm. While SFT endows models with extensive chemical knowledge, it primarily encourages pattern memorization, limiting their ability to perform complex, multi-step reasoning. Crucially, our framework is the only one to generate a chemically-grounded, natural language explanation for its predictions.
Method Baseline Collab. $\Delta$ Interpretability
Template-based
LocalRetro 63.9 66.1 +2.2
GLN 64.2 66.5 +2.3
Semi-template
GraphRetro 63.9 66.2 +2.4
RetroPrime 64.8 67.4 +2.6
Graph2Edits 67.2 70.3 +3.1
Template-free
Retroformer 63.5 65.5 +2.0
UAlign 66.2 69.0 +2.8
Table 2: Top-1 accuracy of individual retrosynthetic specialized models and their collaboration with our Retro-Expert on the USPTO-50K dataset. $\Delta$ denotes the improvement when Retro-Expert collaborates with the specialized model.
@@
Synergistic Enhancement of Specialized Models. Beyond outperforming LLMs, Retro-Expert demonstrates strong synergy when collaborating with different specialized, non-LLM models (Table 2). Retro-Expert surpasses the standalone Top-1 accuracy of these expert models, with performance gains scaling proportionally as the baseline model’s own accuracy increases. These results validate that Retro-Expert effectively reasons over the provided candidate information to optimize its final decision. This collaborative mechanism serves a dual purpose: it boosts the predictive performance of the entire system while making the final choice fully interpretable.
Emergent Capability for Self-reflection and Reasoning. A particularly compelling finding is the emergent capability of Retro-Expert for self-reflection and reasoning, which allows the LLM to overcome the limitations of the collaborating specialized models. The framework does not uncritically accept the candidates provided by specialized models. Instead, it exhibits a capacity for critical analysis and self-correction. As illustrated in Figure 19, when a specialized model fails to provide valid reactant candidates (e.g., they are entirely incorrect), Retro-Expert leverages the available chemical context to autonomously generate a novel and correct prediction. A statistical analysis of such failure cases reveals that this self-reflection mechanism achieved a remarkable 46.2% success rate, demonstrating that it is not a selection model but a generative model that can create new answers.
Decision Space Reactants Top-1 Accuracy (%)
Type Center Reactant
$\checkmark$ 28.9
$\checkmark$ $\checkmark$ 31.0
$\checkmark$ $\checkmark$ 53.7
$\checkmark$ $\checkmark$ $\checkmark$ 66.2
Table 3: Performance of different combinations of reasoning strategies across different subtasks. Sub-tasks with both unchecked indicate the sub-task is skipped and directly execute retrosynthesis prediction.
@@
Retro-Expert 4.17 3.69 4.23
Alation Study
Effective Reasoning Hinges on the Chemical Decision Space. To validate the necessity of constructing chemical decision space, we train LLMs using RL under different decision spaces. As shown in Table 3, When only three candidate reactants are provided, the model’s accuracy is extremely low (28.9%). When gradually add “type”, “center” into the decision space, the performance gradually increase to 66.2%. This indicates that a structured, multi-dimensional decision space is essential to properly ground the LLM’s reasoning and achieve high accuracy.
Model Top-1 Acc (%)
Qwen2.5-7B-Instruct (Yang et al. 2025) 29.33
ChemLLM (Zhang et al. 2024) 15.33
DeepSeek-R1 (Guo et al. 2025) 32.54
Retro-Expert (ours) 57.00
Table 5: Top-1 accuracy (%) of different models on the out-of-distribution ChemBench benchmark.
Interpretability Evaluation
Since explanatory capabilities are largely absent in specialized chemical models, we benchmarked Retro-Expert against general LLMs on the USPTO-50K test set. These baseline models were explicitly prompted to generate a step-by-step rationale. The evaluation combines automatic scoring using GPT-4o with expert human assessment. For the human evaluation, we recruited three independent evaluators, each with experience in synthetic organic chemistry.
@@
Retro-Expert Delivers Credible Reasoning Superior to General LLMs. As detailed in Table 4, Retro-Expert achieves scores near or above 4.0 in the automated GPT-4o assessment, significantly outperforming the general LLM baselines. This strong performance was largely corroborated by human experts, who awarded high scores for Mechanism Accuracy (MA) and Logical Consistency (LC), confirming that Retro-Expert generates clear, coherent, and mechanistically plausible reasoning pathways. Interestingly, our analysis revealed a notable discrepancy in the Factual Correctness (FC) metric, where human experts assigned a more moderate score (3.69) than the automated assessment. This discrepancy reveals that while the model’s reasoning process is robust, it can occasionally make minor factual errors. These demonstrated capabilities are crucial for enhancing the transparency and interpretability of the prediction process, fostering greater trust in its real-world application.
Generalization on Out-of-Distribution Data
We evaluated Retro-Expert’s generalization on the ChemBench out-of-distribution (OOD) benchmark, whose 300 reaction pathways are 94.66% absent from the USPTO-50K dataset, thus reflecting a real-world discovery scenario. To ensure a rigorous and fair comparison, we formulated the identical decision space for all models as a four-option choice task (one ground truth, three distractors).
As shown in Table 5, a stark performance gap emerges, highlighting the critical difference between general reasoning and specialized, principle-driven reasoning. Despite operating within the same chemical decision space, general LLMs struggle on these OOD samples, achieving accuracies around 30%. Their failure stems from an inability to consistently apply chemical principles to navigate the decision space, often defaulting to superficial pattern matching which is ineffective for novel reactions. In contrast, Retro-Expert’s accuracy of 57.00% nearly doubles the baselines. This superior generalization results directly from its training: the Knowledge-Guided Policy Optimization (KGPO) ingrains a transferable, chemistry-principled reasoning policy. This policy enables Retro-Expert to strategically navigate the decision space for novel reactions, demonstrating a crucial step towards reliable AI in chemical discovery.
Case Study: Making the Reasoning Transparent
As visualized on the right side of Figure 2, Retro-Expert generates a human-readable rationale that articulates its step-by-step analysis for a specific molecule. The process begins with the construction of the decision space. For this example, specialized models identify “functional group addition” as a probable reaction type and pinpoint positions 11 and 12 as potential reaction centers. The model’s rationale then demonstrates its navigation of this space. For instance, it correctly discards a candidate reactant, explicitly stating it is ruled out due to the “absence of a clear bromomethyl group”. The logical and verifiable narrative is essential for enhancing an chemical expert’s trust and boosting the model’s practical applicability.
Wet Lab Experiments
@@
To validate Retro-Expert’s practical utility, we extended beyond in-silico (dry-lab) tests to chemical web-lab experiments, demonstrating its capacity as a genuine engine for chemical discovery. We provide direct experimental evidence that Retro-Expert successfully pioneers the first reported synthesis of a molecule that previously lacked any documented production path, and discovers a novel reaction route for an existing compound.
@@
Our experiments successfully validate these two distinct types of chemical discovery. (1) Our Retro-Expert model predicted a new route of Suzuki-Miyaura coupling to synthesize 3-(2-ethoxyphenyl)thiophene (CCOC1=CC=CC=C1C1=CSC=C1) via a reaction between 1-bromo-2-ethoxybenzene (CCOc1ccccc1Br) and thiophen-3-ylboronic acid (OB(O)c1ccsc1). Crucially, verification using the CAS SciFinder database confirmed that while the compound’s structure was known, no synthesis pathway had ever been published for it. Our work thus constitutes its first-ever reported synthesis, which we successfully achieved with a percent yield of 79.3%. (2) For the well-established molecule 1-(4-ethoxyphenyl)ethanone (CCOC1=CC=C(C=C1)C(C)=O), our model identified a novel Jones Oxidation pathway. The target compound can be synthesized via 1-(4-ethoxyphenyl)ethanol (CCOc1ccc(C(C)O)cc1) using chromium trioxide in acidic conditions. This route, confirmed to be an undocumented alternative to existing methods, was also successfully executed in our chemical lab with a yield of 58.82%. These outcomes offer compelling proof that Retro-Expert is a transformative tool for practical chemical discovery.
Conclusion

### v2

@@ 3.5 Knowledge-Guided Policy Optimization
In this objective function, $\pi_{\theta}$ represents the policy model being optimized, while $\pi_{ref}$ is a reference model used to regularize the policy update, with their divergence measured by $\mathbb{D}_{kl}$. The term $q$ denotes a query sampled from the dataset, comprising the candidate set $\mathcal{T}$ and any external domain knowledge $L$. The output $y$ generated by the policy model includes both the complete reasoning path $\mathcal{T}_{LLM}$ and the final predicted reactants $a$. The optimization aims to maximize the expected reward $r_{\phi}(q,y)$. To optimize this objective, we builds upon the Group Relative Policy Optimization (GRPO) algorithm [28]. However, GRPO focus the reward signals solely on the correctness of the final reactant, which limits the model’s performance in chemical reasoning tasks. In practice, a chemist’s trust is earned not by a single, opaque prediction, but by an interpretable reasoning process that aligns with established chemical principles.
To address this limitation, we aim to guide the model toward being “right for the right reasons”, meaning it reaches correct conclusions through chemically valid reasoning. Based on this principle, we propose the Knowledge-Guided Policy Optimization (KGPO) mechanism. The core objective of KGPO is to incentivize the LLM to learn a chemically-grounded reasoning policy that prioritizes the logical validity of the entire pathway, not merely the correctness of the final prediction. It design a novel rule-based, multi-stage reward mechanism tailored to encourage the model to effectively leverage chemical knowledge in the decision space. Specifically, the KGPO framework utilizes a composite reward function that evaluates both the process and outcome. This function is specifically designed to leverage domain knowledge from specialized models, guiding the LLM’s reasoning toward chemical validity while maintaining high accuracy in the final prediction. For a sampled decision path $y$, it is defined as:
$r(\mathcal{T}_{\text{LLM}},y)=\alpha_{\text{1}}\sum_{i=1}^{s}r_{i}+\alpha_{\text{2}}r_{\text{reactant}}+\alpha_{\text{3}}r_{\text{format}},$ (5)
where the term $r_{i}$ represents the binary stage reward for each of the $s$ sub-tasks. It is assigned a value of 1 if the reasoning process contains correct information for the particular sub-task, and 0 otherwise. $r_{\text{reactant}}$ is the reward for the correctness of the final predicted reactants and $r_{\text{format}}$ aim to ensure the required output format.
@@
4 Experiments
4.1 Experimental Setup
@@
Dataset. We conduct experiments using two benchmark dataset: USPTO-50K [24] and USPTO-FULL [6], which contain 50k and 1 million atom-mapped reaction records, respectively. We adopt the same training/validation/test splits (8:1:1) as prior works [6]. Following previous methods [30], we canonicalize the product SMILES and reassign the atom-mapping to the corresponding reactant SMILES based on the canonical ordering.
@@
Baselines. We compare Retro-Expert against various strong baselines, categorized into two primary classes: (1) non-LLM-based approaches, including template-based (LocalRetro [4], GLN [6]), semi-template-based (GraphRetro [30], RetroPrime [37], Graph2Edits [47]), and template-free (Retroformer [34], UAlign [43]) models; (2) LLM-based methods, including chemistry-specialized LLMs and general LLMs.
@@
Evaluation Metrics. Inspired by [45], we adopt two complementary sets of metrics to evaluate the performance between the predicted and golden standardized SMILES. The first set focuses on the textual similarity, using Top-1 accuracy, BLEU, and levenshtein distance. The second set evaluates the chemical similarity, encompassing the validity of the generated SMILES and fingerprint tanimoto similarity (i.e., MACCS, RDK, Morgan). Additionally, we compare the performance of Retro-Expert against retrosynthesis small models using Top-K accuracy.
@@
Implementation Details. In training, we employ two specialized model for space construction, i.e., T5Chem [20] for reaction type prediction, and GraphRetro [30] for reaction center localization and reactant prediction. During inference, ANY retrosynthetic models can be used to provide reactant candidates in a plug-and-play way, without retraining of the LLM. To balance accuracy and optimization efficiency, we use Top-4 candidate predictions from each model. We utilize Qwen2.5-7B-Instruct [41] as the LLM and train it using 9k and 80k samples from USPTO-50K [24] and USPTO-FULL [6] datasets, respectively. During training, to prevent reward hacking, we distribute the correct predictions from the specialized model among the Top-4 candidate positions in a 5:3:2 ratio. This strategy prevents the model from developing a positional bias (e.g., always selecting the first candidate) and forces it to learn the intrinsic quality of predictions. The reward weighting coefficients are set as $\alpha_{1}=1.5$, $\alpha_{2}=1.0$, and $\alpha_{3}=0.2$.
Table 2: Comparison with non-LLM-based models on USPTO-50K [24]. Ours indicates the same LLM that reasons over a decision space constructed by the corresponding retrosynthesis model.
Method Top-K accuracy (%) Interpre- table
1 3 5 10
Retroformer [34] 63.5 82.1 86.3 90.2
LocalRetro [4] 63.9 86.8 92.4 96.3
GraphRetro [30] 63.9 81.5 85.2 88.1
GLN [6] 64.2 79.1 85.2 90.0
UAlign [43] 66.2 86.9 91.9 95.1
Graph2Edits [47] 67.2 87.5 91.5 93.8
Ours (Retroformer) 65.5 83.2 86.6 90.1
Ours (LocalRetro) 66.1 88.9 93.1 96.7
Ours (GraphRetro) 66.2 82.4 85.4 88.1
Ours (GLN) 66.5 79.8 85.3 90.0
Ours (UAlign) 69.0 88.9 92.5 95.4
Ours (Graph2Edits) 70.3 89.4 92.1 94.0
4.2 Comparison with State of the Art
Superiority over LLM-based Approaches. As detailed in Table 1, Retro-Expert significantly outperforms both chemistry-focused foundation models and general LLMs, achieving a Top-1 Accuracy improvement of over 26.64%. This demonstrates that our knowledge-guided reinforcement learning approach effectively enhances the model’s reasoning capability. In contrast, the chemical LLMs trained via SFT tend to replicate memorized reaction patterns for prediction. This strategy exhibits suboptimal performance for retrosynthesis prediction.
Superiority over Non-LLM-based Approaches. Beyond outperforming LLMs, Retro-Expert also demonstrates clear advantages over non-LLM-based models (Table 2). When using different retrosynthesis models to construct the reactant-related decision space, Retro-Expert achieves higher Top-K accuracy than the underlying small models. Moreover, our methods exhibits model-agnostic compatibility, enabling seamless integration with various specialized models at no cost (no retraining required). The performance gains scaling proportionally as the baseline model’s own accuracy increases. These results validate that Retro-Expert effectively reasons over the provided candidate information to optimize its final decision. This collaborative mechanism serves a dual purpose: it boosts the predictive performance of the entire system while making the final choice fully interpretable.
Table 3: Comparison of different retrosynthesis methods on the USPTO-FULL [6] dataset.
Method Top-1 Acc Top-3 Acc
LocalRetro [4] 39.1 53.3
GLN [6] 39.3
Graph2Edits [47] 44.0 60.9
GTA [27] 46.6
EditRetro [13] 47.1 61.2
Ours 49.3 63.0
@@ 4.2 Comparison with State of the Art
Beyond the USPTO-50K dataset, we further evaluate the performance on the larger and more diverse USPTO-FULL dataset. As shown in Table 3, our method achieves superior performance to existing approaches. This demonstrates that our method is capable of handling more complex and and diverse reaction scenarios.
Emergent Capability for Self-reflection. A particularly compelling finding is the emergent capability of Retro-Expert for self-reflection and reasoning, which allows the LLM to overcome the limitations of the input specialized models. The framework does not uncritically accept the candidates provided by specialized models. Instead, it exhibits a capacity for critical analysis and self-correction. As illustrated in Figure 10, when a specialized model fails to provide valid reactant candidates (e.g., they are entirely incorrect or no candidate), Retro-Expert may generate a new correct answer. A statistical analysis of such cases reveals that this self-reflection mechanism achieved a remarkable 40.6% success rate, showing that it is not a selection model but a generative model that can create new answers.
Superior Generalization on Out-of-Distribution Data. We evaluate Retro-Expert’s generalization on the out-of-distribution (OOD) ChemBench benchmark [44], where 94.66% of the 300 reaction pathways are absent from USPTO-50K, reflecting real-world discovery conditions. To ensure fairness, all models shared an identical decision space formulated as a four-option choice task (one ground truth and three distractors). As shown in Table 4, general LLMs achieve accuracy around 30%. In contrast, Retro-Expert attains 57.00% accuracy. This improved generalization is enabled by KGPO, which ingrains a transferable, chemistry-principled reasoning policy.
Table 4: Top-1 accuracy (%) of different models on the out-of-distribution ChemBench benchmark [44].
Method Top-1 Accuracy (%)
ChemLLM [44] 15.33
Qwen2.5-7B-Instruct [41] 29.33
DeepSeek-R1 [11] 32.54
Ours 57.00
@@
Ours 4.17 3.69 4.23
4.3 Interpretability Evaluation
Since explanatory capabilities are largely absent in specialized chemical models, we benchmarked Retro-Expert against general LLMs on the USPTO-50K test set. These baseline models were explicitly prompted to generate a step-by-step rationale. The evaluation combines GPT-4o automatic scoring with expert human assessment. For the human evaluation, we recruited three independent evaluators, each with experience in synthetic organic chemistry.
@@ 4.3 Interpretability Evaluation
Retro-Expert Delivers Credible Reasoning Superior to General LLMs. As detailed in Table 5, Retro-Expert achieves scores near or above 4.0 in the automated GPT-4o assessment, significantly outperforming the general LLM baselines. This strong performance was largely corroborated by human experts, who awarded high scores for Mechanism Accuracy (MA) and Logical Consistency (LC), confirming that Retro-Expert generates clear, coherent, and mechanistically plausible reasoning pathways. These demonstrated capabilities are crucial for enhancing the transparency and interpretability of the prediction process, fostering greater trust in its real-world application.
Table 6: Performance comparison under different decision space.
Decision Space Reactants Top-1 Accuracy (%)
Type Center Reactant
$\checkmark$ 29.1
$\checkmark$ $\checkmark$ 32.6
$\checkmark$ $\checkmark$ 56.7
$\checkmark$ $\checkmark$ $\checkmark$ 70.3
Table 7: Performance comparison using different training policy. Top-1 accuracy quantifies reasoning correctness, and MA, FC, and LC assess the reasoning process.
Setting Reasoning Result Reasoning Process
Top-1 Acc (%) MA FC LC
SFT 48.6 2.75 2.12 2.84
RL (GRPO [28]) 66.8 3.37 3.65 3.49
RL (KGPO) 70.3 4.33 3.97 4.25
4.4 Ablation Study
We conducted a series of ablation studies to validate the core architectural choices of Retro-Expert. Additional studies are provided in the appendix.
Effective Reasoning Hinges on the Chemical Decision Space. To demonstrate the necessity and effectiveness of constructing chemical decision space, we train LLMs using KGPO under different decision space. As shown in Table 6, when only four candidate reactants are provided, the model’s prediction accuracy is extremely low (29.1%). When gradually add “type”, “center” into the decision space, the performance gradually increases to 70.3%. This indicates that a structured, multi-dimensional decision space is essential to properly ground the LLM’s logical reasoning and achieve high accuracy.
Effectiveness of KGPO for Superior Reasoning. To understand the effectiveness of KGPO in enhancing LLM reasoning, we train LLMs using different training strategies. As shown in Table 7, the KGPO-trained LLM achieves superior performance compared to other baselines. SFT encourages pattern memorization for prediction. This limits the model’s ability to flexibly perform logical reasoning, thereby creating performance bottlenecks in reasoning-intensive tasks such as retrosynthesis. While GRPO improves accuracy, it solely focuses on the final answer’s correctness. This may lead the model to exploit logically flawed or chemically implausible shortcuts to arrive at the correct conclusion, resulting in suboptimal reasoning outcomes and paths. In contrast, our KGPO strategy effectively guides the model in utilizing the structured decision space, resulting in more accurate prediction. Furthermore, the strategy notably enhances the quality of the reasoning process, improving both the coherence and accuracy of the generated content.
Figure 4: Visualization of the novel reaction pathways proposed by Retro-Expert and validated through wet-lab experiments. The details are provided in the appendix.
4.5 Wet Lab Experiments
@@
To validate Retro-Expert’s practical utility, we extended beyond in-silico (dry-lab) tests to chemical web-lab experiments, demonstrating its capacity as a genuine engine for chemical discovery. We provide direct experimental evidence that Retro-Expert can successfully propose a feasible synthesis of a molecule that previously lacked any documented production path, and discovers a novel reaction route for an existing compound.
@@
Our experiments successfully validate these two distinct types of chemical discovery. (1) Our Retro-Expert model predicted a new route of Suzuki-Miyaura coupling to synthesize 3-(2-ethoxyphenyl)thiophene (CCOC1=CC=CC=C1C1=CSC=C1) via a reaction between 1-bromo-2-ethoxybenzene (CCOc1ccccc1Br) and thiophen-3-ylboronic acid (OB(O)c1ccsc1). Crucially, verification using the CAS SciFinder database confirmed that while the compound’s structure was known, no synthesis pathway had ever been published for it. Our work thus constitutes its first-ever reported synthesis, which we successfully achieved with a percent yield of 79.3%. (2) For the well-established molecule 1-(4-ethoxyphenyl)ethanone (CCOC1=CC=C(C=C1)C(C)=O), our model identified a novel Jones Oxidation pathway. The target compound can be synthesized via 1-(4-ethoxyphenyl)ethanol (CCOc1ccc(C(C)O)cc1) using chromium trioxide in acidic conditions. This route, confirmed to be an undocumented alternative to existing methods, was also successfully executed in our chemical lab with a yield of 58.82%. These outcomes offer compelling proof that Retro-Expert is a effective tool for practical chemical discovery.
Table 8: Comparison of different methods on multi-step retrosynthesis (the test set of PaRoutes [10] dataset).
Method Top-1 Acc (%)
DFPN [17] 11
Retro [3] 15
MCTS [26] 17
DMS-Deep [29] 36
Ours 38
4.6 Case Study: Making Reasoning Transparent
As visualized on the right side of Figure 2, Retro-Expert generates a human-readable rationale that articulates its critical analysis for a specific molecule. The process begins with the construction of the decision space. For this example, specialized models identify “functional group addition” as a probable reaction type and pinpoint positions 11 and 12 as potential reaction centers. The model’s rationale then demonstrates its navigation of this space. For instance, it correctly discards a candidate reactant, explicitly stating that it is ruled out due to the “absence of a clear bromomethyl group”. The logical and verifiable narrative is essential for enhancing a chemical expert’s trust and boosting the model’s practical applicability.
4.7 Multi-Step Retrosynthesis Prediction
We extend Retro-Expert to enable multi-step retrosynthetic prediction based on the PaRoutes dataset [10]. Specifically, we employ the DMS-Deep [29] model to generate candidates that form the decision space. As shown in Table 8, we report the Top-1 accuracy of different models on the test set. Our proposed method achieves the highest prediction accuracy while generating a natural language-based reasoning process. These results underscore the potential of our model for practical retrosynthesis predictions.
5 Conclusion